Simba Company’s standard materials cost per unit of output is $10.12 (2.30 pounds x $4.40). During July, the company purchases and uses 2,898 pounds of materials costing $15,359 in making 1,400 units of finished product. Compute the total, price, and quantity materials variances. (Round per unit values to 2 decimal places, e.g. 52.75 and final answers to 0 decimal places, e.g. 52.)

Total materials variance

Materials price variance

Materials quantity variance
Answer #1
Budgeted Data
Material used per output = 2.3 pounds
Price per pound of material = $4.40
Actual Data
Material used per output = 2898 pounds/1400 units = 2.07 pounds
Price per pound of material = $15359/2898 pounds = $5.30
Total Material Variance = (Budgeted material used * budgeted price * 1400 UNITS) - (Actual material used *actual price * 1400 UNITS)
= (2.3*4.4*1400)-(2.07*5.30*1400)
= 14168-15359
= $1,191 (Unfavourable)
Material price variance = (budgeted price - Actual Price) * Budgeted material used per unit of output * 1400 units
= (4.40-5.30)*2.3*1400
= $2,898 (Unfavourable)
Material quantity variance = (budgeted quantity per unit - Actual quantity per unit) * Actual price * 1400 units
= (2.3 -2.07) * 5.30 * 1400
= $1,707 (Favourable)
Proof:
Total Material Variance = Material price variance + Material quantity variance
($1,191) ($2,898) + $1,707
($1,191)
Hence proved!
Answer #2
Materials Price Variance =Actual Quantity x (Actual Price - Standard Price)
=2898 x ($5.30 - $4.40)
=$2,608 UUnfavorable



Materials Quantity Variance =Standard Price x (Actual Quantity used - Standard Quantity allowed)
=$4.40 x (2898 - 3220*)
=$1,417 FUnfavorable



Total Materials Variance =$2,608U + $1,417F
=$ 1,191 U

*1400x2.30=3220

answered by: Dang Tuan Hiep